John is one fifth of his father’s age. If his dad was 32 when John was born, how old are they both now?
Answer by dkppathak(439)

let age of father is 5x yrs
and age of john is x yrs
age of father when john born was 32 yrs
present age of father when john age is x +32 yrs with consideration of both conditions we can say
5x=x+32
5x-x=32
4x=32
x=32/4
x=8yrs
john present age is 8 yrs
and father present age is 40 yrs
answer father 40 yrs
john age is 8yrs